Exercise 5-17 (Part Level Submission) Pharoah Bucket Co., a manufacturer of rain barrels, had the following data for 2019. Sa(b) Your answer is correct. What is the break-even point in dollars? Break-even point 574751(c) Your answer is correct. What is the margin of safety in dollars and as a ratio? Margin of safety 57475 $ Margin of safety(d) If the company wishes to increase its total dollar contribution margin by 30% in 2020, by how much will it need to increa
Answer #1

Solution d:

Existing total contribution margin = ($55 - $33) * 2090 = $45,980

Required contribution margin = $45,980 * 130% = $59,774

Required sales to earn desired contribution margin = $59,774 / 40% = $149,435

Total increase in sales required = $149,435 - (2090*$55) = $34,485